OpenStudy (anonymous):

write the equation of a parabola with the shape as f(x)=-2x^ but with the vertex (-3,2)???

OpenStudy (cwrw238):

y = -2(x + 3)^2 + 2

OpenStudy (anonymous):

y-2=-2(x-(-3))^2

OpenStudy (anonymous):

formula is(y-k)=-2(x-h)^2 where (h,k) is vertex

OpenStudy (cwrw238):

rearranging y = -2x^2 - 12x - 16